### v2

@@ 3 Problem Definition
where $x(k)\in\mathbb{R}^{n_{x}}$ is the state, $u(k)\in\mathbb{R}^{n_{u}}$ is a known input, and $y_{i}(k)\in\mathbb{R}^{m_{i}}$ is the measured output of the $i$-th sensor with $i\in\mathbb{Z}_{[1,p]}$. The vector $w(k)\in\mathcal{W}$ represents the process noise, which is bounded and assumed to be contained in the zonotope $\mathcal{W}=\langle c_{w},G_{w}\rangle$, and the vector $v_{i}(k)\in\mathcal{V}_{i}$ represents the measurement noise of the $i$-th sensor, which is also bounded and assumed to be contained in the zonotope $\mathcal{V}_{i}=\langle c_{v_{i}},G_{v_{i}}\rangle$, for every $i\in\mathbb{Z}_{[1,p]}$. Sensor $i$’s measurement at time $k$ may be corrupted by an arbitrary and unbounded attack signal $a_{i}(k)\in\mathbb{R}^{m_{i}}$, which is assumed to be injected by a malicious attacker.
@@
    Maximum number of attacked sensors: The attacker can attack up to $q\in\mathbb{Z}_{[0,p-1]}$ number of sensors. Although the upper bound $q$ is known, the exact number and/or the set of attacked sensors need not be known.
@@
    Redundant observability: There exists $c_{\mathsf{J}}\leq p-q$ such that, for every subset $\mathsf{J}\subset\mathbb{Z}_{[1,p]}$ of sensors with cardinality $|\mathsf{J}|=c_{\mathsf{J}}$, the pair $(A,C_{\mathsf{J}})$ is observable, where $C_{\mathsf{J}}$ is obtained by stacking all $C_{i}$, $i\in\mathsf{J}$, in row blocks.
@@
    Knowledge of an initial set: At time $k=0$, we know a compact set $\mathcal{X}_{0}\subset\mathbb{R}^{n_{x}}$ containing the initial state $x(0)\in\mathcal{X}_{0}$.
@@
Assumption 1(i) is fundamental in this paper because it ensures that, at every time $k\in\mathbb{Z}_{\geq 0}$, there exists a set of uncompromised (or safe) sensors:
@@
such that $a_{i}(k)=0_{m_{i}}$ for every $i\in\mathsf{S}_{k}$. This, along with the redundant observability (Assumption 1(ii)), allows us to ensure that the true state can be theoretically reconstructed from the set of uncompromised sensors under the absence of noise. In addition, the assumption entails that the attacker, even though omniscient about the system dynamics and noise bounds, has limited resources at hand and cannot attack all the sensors. We remark that this assumption is not restrictive because it neither restricts the set of attacked sensors to be static with respect to time nor requires that $q$ is less than half the number of sensors $p$ — an assumption that is fundamental in the point-based secure state estimation literature. On the contrary, at any time instant, our problem setup allows the attacker to inject arbitrary signals to any subset of sensors with cardinality less than or equal to $q$, where $q$ is only required to be strictly less than $p$ — meaning that at least one sensor needs to be safe at any time instant.
Assumption 1(ii) is required to enable decentralized set-based operations for secure state estimation without violating the robustness guarantees. Moreover, because Assumption 1(i) allows the attacker to attack up to $q\leq p-1$ sensors, it is necessary that the observability is guaranteed from the remaining safe sensors.
Assumption 1(iii) is required to initialize the set-based state estimation algorithm with $\hat{\mathcal{X}}_{0}=\mathcal{X}_{0}$ such that the inclusion of the state $x(k)$ in the estimated set $\hat{\mathcal{X}}_{k}\subset\mathbb{R}^{n_{x}}$ can be guaranteed. Note that Assumption (iii) does not require any bound on the radius of $\mathcal{X}_{0}$. The only requirement is that the initial set $\mathcal{X}_{0}$ contains $x(0)$. Finally, Assumption 1(iv) is a standard assumption in robust estimation and control literature; see [42] and [15].
@@
  1. P1.
    Estimate a set ${\hat{\mathcal{X}}_{k}\subset\mathcal{X}}$ guaranteeing the inclusion $x(k)\in\hat{\mathcal{X}}_{k}$ for every $k\in\mathbb{Z}_{\geq 1}$, where $x(k)$ is the true state of (3).
  2. P2.
    Provide conditions on the magnitude of the attack signals $a_{i}(k)$ such that they can be detected, identified, and/or filtered.
  3. P3.
    Under what conditions, $\hat{\mathcal{X}}_{k}$ satisfies the following stability condition
    $\mathrm{rad}(\hat{\mathcal{X}}_{k})\leq\beta(\mathrm{rad}(\mathcal{X}),k)+\gamma_{1}(\mathrm{rad}(\mathcal{W}))+\gamma_{2}(\mathrm{rad}(\mathcal{V}))$ (4)
    irrespective of the attack signals $a_{i}(k)$, where $\beta$ is a class-$\mathcal{KL}$ function, $\gamma_{1},\gamma_{2}$ are class-$\mathcal{K}_{\infty}$ function [^2] A continuous function $\gamma:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ is a class $\mathcal{K}$ function, if it is strictly increasing and $\gamma(0)=0$. It is class $\mathcal{K}_{\infty}$ if it is also unbounded. A continuous function $\beta:\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ is a class $\mathcal{KL}$ function, if: (i) $\beta(.,s)$ is a class $\mathcal{K}$ function for any $s\geq 0$; (ii) $\beta(r,.)$ is non-increasing and (iii) $\beta(r,s)\to 0$ as $s\to\infty$ for any $r\geq 0$., and $\mathrm{rad}(\mathcal{V})=\max(\mathrm{rad}(\mathcal{V}_{1}),\dots,\mathrm{rad}(\mathcal{V}_{p}))$.
@@ 4 Secure Set-based State Estimation Algorithm
This section presents our secure set-based state estimation (S3E) algorithm. The algorithm comprises four steps that are summarized in Algorithm 1. In the following subsections, we describe each step of the algorithm in detail, provide a comparison with point-based estimators, and address the complexity issue.
@@ 4.1 Time Update
$\hat{\mathcal{X}}_{k|k-1}=A\hat{\mathcal{X}}_{k-1}\oplus Bu(k-1)\oplus\mathcal{W},$ (5)
@@
Given $x(k-1)\in\hat{\mathcal{X}}_{k-1}$, the time update set $\hat{\mathcal{X}}_{k|k-1}\ni x(k)$ is the set of states to which the system can evolve at time $k\in\mathbb{Z}_{\geq 0}$ subject to the model $(A,B)$, the input $u(k-1)$, and the noise zonotope $\mathcal{W}$. Although the attacker cannot directly influence the time update set as seen from (5), it can do so indirectly through the previous estimate $\hat{\mathcal{X}}_{k-1}$. Thus, the resilience against sensor attacks is achieved by carefully devising the estimated set $\hat{\mathcal{X}}_{k}$.
@@ 4.2 Measurement Update
be sensor $i$’s output measurement set. Then, the measurement update of sensor $i$ is given by the following generalized intersection of the time update set $\hat{\mathcal{X}}_{k|k-1}$ with the output measurement set $\mathcal{Y}_{k}^{i}$ of sensor $i$ as stated below
@@
The measurement update $\hat{\mathcal{Z}}_{k}^{i}$ is a subset of the time update set $\hat{\mathcal{X}}_{k|k-1}$ that is consistent with the measurement of sensor $i$. That is, given the measurement noise zonotope $\mathcal{V}_{i}$, $\hat{\mathcal{Z}}_{k}^{i}$ is the set of states $x(k)\in\hat{\mathcal{X}}_{k|k-1}$ that could have produced the measurement $y_{i}(k)$. Notice that $\hat{\mathcal{Z}}_{k}^{i}$ could be an empty set if the magnitude of the attack signal $a_{i}(k)$ is sufficiently large.
@@ 4.3 Set-based Agreement Protocol
Let $n_{\mathsf{J}}:={p\choose c_{\mathsf{J}}}$, where $c_{\mathsf{J}}\leq p-q$ is the redundant observability parameter. Let $\mathsf{J}_{1},\dots,\mathsf{J}_{n_{\mathsf{J}}}\subset\mathbb{Z}_{[1,p]}$ be the disjoint subsets of size $|\mathsf{J}_{h}|=c_{\mathsf{J}}$, for $h\in\mathbb{Z}_{[1,n_{\mathsf{J}}]}$. The agreement protocol between different measurement updates is given by
@@
We first show that the measurement update $\hat{\mathcal{Z}}_{k}^{i}$ contains the true state $x(k)$ if sensor $i$ is in the uncompromised set of sensors $\mathsf{S}_{k}$ at time $k$. If $i$ is uncompromised at time $k$, then $a_{i}(k)=0$ and $y_{i}(k)=C_{i}x(k)+v_{i}(k)$. From (6), we have $\mathcal{Y}_{k}^{i}=\langle c_{y_{i}(k)},G_{v_{i}}\rangle$ with
$c_{y_{i}(k)}=y_{i}(k)-c_{v_{i}}=C_{i}x(k)+v_{i}(k)-c_{v_{i}}.$
As $\mathcal{Y}_{k}^{i}$ and the noise zonotope $\mathcal{V}_{i}$ have the same generator matrix $G_{v_{i}}$, we have $\mathrm{rad}(\mathcal{Y}_{k}^{i})=\mathrm{rad}(\mathcal{V}_{i})$. Moreover, it holds that
@@
Thus,
$c_{y_{i}(k)}-v_{i}(k)+c_{v_{i}}=C_{i}x(k)\in\mathcal{Y}_{k}^{i}$
because $\|v_{i}(k)-c_{v_{i}}\|_{\infty}\leq\mathrm{rad}(\mathcal{V}_{i})=\mathrm{rad}(\mathcal{Y}_{k}^{i})$. Since $x(k)\in\hat{\mathcal{X}}_{k|k-1}$ and $C_{i}x(k)\in\mathcal{Y}_{k}^{i}$, it implies that $\hat{\mathcal{Z}}_{k}^{i}$ in (7) is non-empty and contains $x(k)$.
We then show that, at every time $k$, there is a subset $\mathsf{J}_{h}$ of sensors with $|\mathsf{J}_{h}|=c_{\mathsf{J}}$ such that $\mathsf{J}_{h}\subseteq\mathsf{S}_{k}$. This is straightforward because $q$ out of $p$ sensors are compromised and $c_{\mathsf{J}}\leq p-q$, so there must be at least one subset $\mathsf{J}_{h}$ with $|\mathsf{J}_{h}|=c_{\mathsf{J}}$ that contains the uncompromised sensors from the guaranteed $p-q$ uncompromised sensors. In other words, at every time $k\in\mathbb{Z}_{\geq 0}$, there exists $h\in\mathbb{Z}_{[1,n_{\mathsf{J}}]}$ such that $\mathsf{J}_{h}\subseteq\mathsf{S}_{k}$. Since $x(k)\in\hat{\mathcal{Z}}_{k}^{i}$ for every $i\in\mathsf{J}_{h}\subseteq\mathsf{S}_{k}$, the intersection in (8) is non-empty and $x(k)\in\mathcal{I}_{k}^{h}$. ∎
@@
It is equivalently true to say that if the agreement set $\mathcal{I}_{k}^{h}$ is empty, then there must be at least one sensor in $\mathsf{J}_{h}$ that is compromised by the attacker. Notice that this is only a sufficient condition for detecting an attack because a stealthy attacker can design an attack that ensures that $\mathcal{I}_{k}^{h}$ is non-empty at any time $k$. However, this comes at a cost to the attacker, which is a discussion reserved for later sections in the paper.
@@ 4.4 Estimated Set
We prove this result by induction. By Assumption 1(iii), we have $x(0)\in\mathcal{X}_{0}=\hat{\mathcal{X}}_{0}$. Because $w(0)\in\mathcal{W}$, the inclusion holds for the time update at $k=1$,
$x(1)\in\hat{\mathcal{X}}_{1|0}=A\mathcal{X}\oplus Bu(0)\oplus\mathcal{W}.$
Therefore, Theorem 1 can be applied, implying the existence of $h\in\mathbb{Z}_{[1,n_{\mathsf{J}}]}$ such that $x(1)\in\mathcal{I}_{1}^{h}$, which gives
$x(1)\in\hat{\mathcal{X}}_{1}=\bigcup_{h\in\mathbb{Z}_{[1,n_{\mathsf{J}}]}}\mathcal{I}_{1}^{h}.$
This, in turn, implies that the time update at $k=2$ contains the true state,
$x(2)\in\hat{\mathcal{X}}_{2|1}=A\hat{\mathcal{X}}_{1}\oplus Bu(1)\oplus\mathcal{W}.$
Now, let us assume $x(k^{\prime})\in\hat{\mathcal{X}}_{k^{\prime}|k^{\prime}-1}$ for some $k^{\prime}\in\mathbb{Z}_{\geq 2}$. Then, by Theorem 1, there exists $h\in\mathbb{Z}_{[1,n_{c}]}$ such that $x(k^{\prime})\in\mathcal{I}_{k^{\prime}}^{h}$, implying
$x(k^{\prime})\in\hat{\mathcal{X}}_{k^{\prime}}=\bigcup_{h\in\mathbb{Z}_{[1,n_{\mathsf{J}}]}}\mathcal{I}_{k^{\prime}}^{h}.$
Therefore, we have the inclusion at the next time update,
$x(k^{\prime}+1)\in\hat{\mathcal{X}}_{k^{\prime}+1|k^{\prime}}=A\hat{\mathcal{X}}_{k^{\prime}}\oplus Bu(k^{\prime})\oplus\mathcal{W}.$
@@
Thus, the proof is completed because we showed that, for every $k\in\mathbb{Z}_{\geq 1}$, $x(k-1)\in\hat{\mathcal{X}}_{k-1|k-2}$ implies $x(k-1)\in\hat{\mathcal{X}}_{k-1}$, which, in turn, implies $x(k)\in\hat{\mathcal{X}}_{k|k-1}$. Hence, $x(k)\in\hat{\mathcal{X}}_{k}$. ∎
@@
Although the above theorem guarantees the inclusion of the true state, it is important to remark that the number of (constrained) zonotopes in the estimated set $\hat{\mathcal{X}}_{k}$ in (9) may increase with respect to time if the attacker is intelligent and stealthy. We address this issue in Section 4.8 by proposing several techniques that facilitate the computational efficiency of the algorithm. However, it is worth mentioning that the proposed estimation algorithm is resilient because the attacker cannot deteriorate the estimation accuracy arbitrarily over time. If a sensor is injected with a large-valued attack signal, it will be automatically discarded by either yielding an empty measurement update or an empty agreement sets.
The overall algorithm is summarized in Algorithm 1.
Algorithm 1 Secure set-based state estimation
Initialize: $\hat{\mathcal{X}}_{0}=\mathcal{X}_{0}$
for $k=1,2,3,\dots$ do
     Time update (5): $\hat{\mathcal{X}}_{k|k-1}$.
     Estimate (9): $\hat{\mathcal{X}}_{k}$.
end for
4.5 Illustrative Example

### v3

@@ 3 Problem Definition
where $x(k)\in\mathbb{R}^{n_{x}}$ is the state, $u(k)\in\mathbb{R}^{n_{u}}$ is a known input, and $y_{i}(k)\in\mathbb{R}^{m_{i}}$ is the measured output of the $i$-th sensor with $i\in\mathbb{Z}_{[1,p]}$. The vector $w(k)\in\mathcal{W}$ represents the process noise, which is assumed to be contained in the zonotope $\mathcal{W}=\langle c_{w},G_{w}\rangle$, and the vector $v_{i}(k)\in\mathcal{V}_{i}$ represents the measurement noise of the $i$-th sensor, which is assumed to be contained in the zonotope $\mathcal{V}_{i}=\langle c_{v_{i}},G_{v_{i}}\rangle$, for every $i\in\mathbb{Z}_{[1,p]}$. Sensor $i$’s measurement at time $k$ may be corrupted by an arbitrary and unbounded attack signal $a_{i}(k)\in\mathbb{R}^{m_{i}}$, which is assumed to be designed by a malicious attacker.
@@
    Maximum number of attacked sensors: The attacker can attack up to $q\leq p-1$ number of sensors. Although the upper bound $q$ is known, the exact number and the set of attacked sensors do not need to be known.
@@
    Redundant observability: There exists $c_{\mathsf{J}}\leq p-q$ such that, for every subset $\mathsf{J}\subset\mathbb{Z}_{[1,p]}$ of sensors with cardinality $|\mathsf{J}|=c_{\mathsf{J}}$, the pair $(A,C_{\mathsf{J}})$ is observable, i.e., $\mathrm{rank}[\begin{array}[]{cccc}C_{\mathsf{J}}^{\top}&(C_{\mathsf{J}}A)^{\top}&\dots&(C_{\mathsf{J}}A^{n_{x}-1})^{\top}\end{array}]^{\top}=n_{x}$, where $C_{\mathsf{J}}$ is obtained by stacking all $C_{i}$, $i\in\mathsf{J}$, in row blocks.
@@
    Knowledge of an initial set: At time $k=0$, we know a bounded set $\mathcal{X}_{0}\subset\mathbb{R}^{n_{x}}$ containing the initial state $x(0)\in\mathcal{X}_{0}$.
@@
Assumption 1(i) restricts the maximum number of attacked sensors at each time instant so that at least one sensor is uncompromised. This assumption is fundamental in this paper, ensuring that, at every time $k\in\mathbb{Z}_{\geq 0}$, there exists a set of uncompromised (or safe) sensors
@@
such that $a_{i}(k)=0_{m_{i}}$ for every $i\in\mathsf{S}_{k}$. This, along with the redundant observability (Assumption 1(ii)), allows us to ensure that the true state can be theoretically reconstructed from the set of uncompromised sensors in the absence of noise. In addition, the assumption entails that the attacker, even though omniscient about the system dynamics and noise bounds, has limited resources at hand and cannot attack all the sensors. We remark that this assumption is not restrictive because it neither restricts the set of attacked sensors to be static with respect to time nor requires that $q$ is less than half the number of sensors $p$, which is the fundamental assumption in the point-based secure state estimation literature. On the contrary, at any time instant, our problem setup allows the attacker to inject arbitrary signals to any subset of sensors with cardinality less than or equal to $q$, where $q$ is only required to be strictly less than $p$.
Assumption 1(iii) is required to initialize the set-based state estimation algorithm with $\hat{\mathcal{X}}_{0}=\mathcal{X}_{0}$ such that the inclusion of the state $x(k)$ in the estimated set $\hat{\mathcal{X}}_{k}\subset\mathbb{R}^{n_{x}}$ can be guaranteed. Note that Assumption (iii) does not require any bound on the radius of $\mathcal{X}_{0}$. Finally, Assumption 1(iv) is a standard assumption in robust estimation and control [44, 15].
Problem Statement
@@
  1. i.
    Estimate a set $\hat{\mathcal{X}}_{k}$ guaranteeing the inclusion of the true state, i.e., $x(k)\in\hat{\mathcal{X}}_{k}$ for every $k\in\mathbb{Z}_{\geq 1}$.
  2. ii.
    Provide conditions such that the attack signals $a_{i}(k)$ can be detected, identified, and/or filtered [^1] These terms are defined later in Section 5..
@@ 4 Secure Set-based State Estimation Algorithm
This section presents the secure set-based state estimation (S3E) algorithm, which is summarized below in Algorithm 1. In the following subsections, we describe all the steps of the algorithm in detail, provide a comparison with the point-based estimators, and present strategies to manage the complexity of the algorithm.
Algorithm 1 Secure set-based state estimation
Initialize: $\hat{\mathcal{X}}_{0}=\mathcal{X}_{0}$
for $k=1,2,3,\dots$ do
  Time update (4): $\hat{\mathcal{X}}_{k|k-1}$
  Estimate (8): $\hat{\mathcal{X}}_{k}$
end for
@@ 4.1 Time Update
$\hat{\mathcal{X}}_{k|k-1}=A\hat{\mathcal{X}}_{k-1}\oplus Bu(k-1)\oplus\mathcal{W}$ (4)
@@
Given $x(k-1)\in\hat{\mathcal{X}}_{k-1}$, the time update set $\hat{\mathcal{X}}_{k|k-1}\ni x(k)$ is the set of states to which the system can evolve at time $k\in\mathbb{Z}_{\geq 0}$ subject to the model $(A,B)$, the input u(k1)u(k-1), and the noise zonotope $\mathcal{W}$. Although the attacker cannot directly influence the time update set as seen from (4), it can do so indirectly through the previous estimate $\hat{\mathcal{X}}_{k-1}$. Thus, the resilience against sensor attacks is achieved by carefully devising the estimated set $\hat{\mathcal{X}}_{k}$.
@@ 4.2 Measurement Update
be sensor $i$’s output measurement set. Then, the measurement update of sensor $i$ is given by the following generalized intersection of the time update set $\hat{\mathcal{X}}_{k|k-1}$ with the output measurement set $\mathcal{Y}_{k}^{i}$ of sensor $i$
@@
The measurement update $\hat{\mathcal{Z}}_{k}^{i}$ is a subset of the time update set $\hat{\mathcal{X}}_{k|k-1}$ that is consistent with the measurement $y_{i}(k)$ of sensor $i$. That is, given the measurement noise zonotope $\mathcal{V}_{i}$, $\hat{\mathcal{Z}}_{k}^{i}$ is the set of states $x(k)\in\hat{\mathcal{X}}_{k|k-1}$ that could have produced the measurement $y_{i}(k)$. Notice that $\hat{\mathcal{Z}}_{k}^{i}$ can be an empty set if the sensor $i$’s measurement is inconsistent with the time update.
@@ 4.3 Set-based Agreement Protocol
Let $n_{\mathsf{J}}\coloneq{p\choose c_{\mathsf{J}}}$, where $c_{\mathsf{J}}\leq p-q$ is the redundant observability parameter. Let $\mathsf{J}_{1},\dots,\mathsf{J}_{n_{\mathsf{J}}}\subset\mathbb{Z}_{[1,p]}$ be the disjoint subsets with $|\mathsf{J}_{h}|=c_{\mathsf{J}}$, for $h\in\mathbb{Z}_{[1,n_{\mathsf{J}}]}$. The agreement protocol between measurement updates in each $\mathsf{J}_{h}$ is given by
@@
We first show that the measurement update $\hat{\mathcal{Z}}_{k}^{i}$ contains the true state $x(k)$ if sensor $i$ is in the uncompromised set of sensors $\mathsf{S}_{k}$ at time $k$. If $i$ is uncompromised at time $k$, then $a_{i}(k)=0$ and $y_{i}(k)=C_{i}x(k)+v_{i}(k)$. From (5), we have $\mathcal{Y}_{k}^{i}=\langle c_{y_{i}(k)},G_{v_{i}}\rangle$ with $c_{y_{i}(k)}=y_{i}(k)-c_{v_{i}}=C_{i}x(k)+v_{i}(k)-c_{v_{i}}.$ As $\mathcal{Y}_{k}^{i}$ and the noise zonotope $\mathcal{V}_{i}$ have the same generator matrix $G_{v_{i}}$, we have $\mathrm{rad}(\mathcal{Y}_{k}^{i})=\mathrm{rad}(\mathcal{V}_{i})$. Moreover, it holds that
@@
Thus, $c_{y_{i}(k)}-v_{i}(k)+c_{v_{i}}=C_{i}x(k)\in\mathcal{Y}_{k}^{i}$ because $\|v_{i}(k)-c_{v_{i}}\|_{\infty}\leq\mathrm{rad}(\mathcal{V}_{i})=\mathrm{rad}(\mathcal{Y}_{k}^{i})$. Since $x(k)\in\hat{\mathcal{X}}_{k|k-1}$ and $C_{i}x(k)\in\mathcal{Y}_{k}^{i}$, it implies that $\hat{\mathcal{Z}}_{k}^{i}$ in (6) is non-empty and contains $x(k)$.
We then show that, at every time $k$, there is a subset $\mathsf{J}_{h}$ of sensors with $|\mathsf{J}_{h}|=c_{\mathsf{J}}$ such that $\mathsf{J}_{h}\subseteq\mathsf{S}_{k}$. This is straightforward because $q$ out of $p$ sensors are compromised and $c_{\mathsf{J}}\leq p-q$, so there must be at least one subset $\mathsf{J}_{h}$ with $|\mathsf{J}_{h}|=c_{\mathsf{J}}$ that contains the uncompromised sensors from the guaranteed pqp-q uncompromised sensors. In other words, at every time $k\in\mathbb{Z}_{\geq 0}$, there exists $h\in\mathbb{Z}_{[1,n_{\mathsf{J}}]}$ such that $\mathsf{J}_{h}\subseteq\mathsf{S}_{k}$. Since $x(k)\in\hat{\mathcal{Z}}_{k}^{i}$ for every $i\in\mathsf{J}_{h}\subseteq\mathsf{S}_{k}$, the intersection in (7) is non-empty and $x(k)\in\mathcal{I}_{k}^{h}$. ∎
@@
It is equivalently true to say that if the agreement set $\mathcal{I}_{k}^{h}$ is empty, then there must be at least one sensor in $\mathsf{J}_{h}$ that is compromised by the attacker. Notice that this is only a sufficient condition for detecting an attack because a stealthy attacker can design an attack that ensures that $\mathcal{I}_{k}^{h}$ is non-empty at any time $k$. However, this comes at a cost to the attacker, which is a discussion reserved for the next section in the paper.
@@ 4.4 Estimated Set
We prove this result by induction. By Assumption 1(iii), we have $x(0)\in\mathcal{X}_{0}=\hat{\mathcal{X}}_{0}$. Because $w(0)\in\mathcal{W}$, the inclusion holds for the time update at $k=1$, $x(1)\in\hat{\mathcal{X}}_{1|0}=A\mathcal{X}\oplus Bu(0)\oplus\mathcal{W}.$ Therefore, Theorem 1 can be applied, implying the existence of $h\in\mathbb{Z}_{[1,n_{\mathsf{J}}]}$ such that $x(1)\in\mathcal{I}_{1}^{h}$, which gives $x(1)\in\hat{\mathcal{X}}_{1}=\bigcup_{h\in\mathbb{Z}_{[1,n_{\mathsf{J}}]}}\mathcal{I}_{1}^{h}.$ This, in turn, implies that the time update at $k=2$ contains the true state, $x(2)\in\hat{\mathcal{X}}_{2|1}=A\hat{\mathcal{X}}_{1}\oplus Bu(1)\oplus\mathcal{W}.$
@@
Now, let us assume $x(k^{\prime})\in\hat{\mathcal{X}}_{k^{\prime}|k^{\prime}-1}$ for some $k^{\prime}\in\mathbb{Z}_{\geq 2}$. Then, by Theorem 1, there exists $h\in\mathbb{Z}_{[1,n_{c}]}$ such that $x(k^{\prime})\in\mathcal{I}_{k^{\prime}}^{h}$, implying $x(k^{\prime})\in\hat{\mathcal{X}}_{k^{\prime}}=\bigcup_{h\in\mathbb{Z}_{[1,n_{\mathsf{J}}]}}\mathcal{I}_{k^{\prime}}^{h}.$ Therefore, we have the inclusion at the next time update, $x(k^{\prime}+1)\in\hat{\mathcal{X}}_{k^{\prime}+1|k^{\prime}}=A\hat{\mathcal{X}}_{k^{\prime}}\oplus Bu(k^{\prime})\oplus\mathcal{W}.$ Thus, the proof is completed because we showed that, for every $k\in\mathbb{Z}_{\geq 1}$, $x(k-1)\in\hat{\mathcal{X}}_{k-1|k-2}$ implies $x(k-1)\in\hat{\mathcal{X}}_{k-1}$, which, in turn, implies $x(k)\in\hat{\mathcal{X}}_{k|k-1}$. Hence, $x(k)\in\hat{\mathcal{X}}_{k}$. ∎
@@
Although the above theorem guarantees the inclusion of the true state, it is important to remark that the number of (constrained) zonotopes in the estimated set $\hat{\mathcal{X}}_{k}$ may increase with respect to time if the attack is stealthy. We address this issue in Section 4.8 by proposing several techniques that manage the computational efficiency of the algorithm. However, it is worth mentioning that the proposed estimation algorithm is resilient because the attacker cannot deteriorate the estimation accuracy arbitrarily over time. If a sensor is injected with a large-valued attack signal, it will be automatically discarded by either yielding an empty measurement update or an empty agreement set (see Section 5).
4.5 Illustrative and numerical examples